OpenStudy (cicilybailey14):

Is -1 a solution of 4y + 3 <= -7?

OpenStudy (ahsome):

Change \(y\) to \(-1\) and see if it works out: \[4y + 3\le -7\]\[4*-1 + 3\le -7\]\[-4 + 3\le -7\]Go from there

OpenStudy (anonymous):

y<=-10/4

OpenStudy (cicilybailey14):

Thx @Ahsome

OpenStudy (ahsome):

No problem @Cicilybailey14 :D